Tyrone is twice as old as his daughter. In six years Tyrone's age will be three times what his daughter's age was six years ago. How old are they at present?
Answer by jainenderkapoor(61)

In this question we have to make equation and solve the equation.
SO how to make the equation in the first step.
We have to find ages of two persons Tyrone and his daughter.
LET us suppose that there present ages are
Tyrone -- x years and Daughter y years
we get the first equations as
x = 2y ------ (1)
Now what will be Tyrone's age after 6 years ? YES (x + 6)
and his daughter's age 6 years ago ? YES (y - 6)
So we get the second equation as
(x + 6) = 3(y - 6) --- (2)

Now putting the value of x from (1) in (2) we get
(2y + 6) = 3(y - 6)
2y + 6 = 3y - 18
Bringing the terms with y on one side and the remaining terms on other side we get
2y - 3y = -18 -6
-y = -24
Cancelling - from both sides we get
y = 24
Putting value of y= 24 in (1) we get
x = 2(24) = 48
So age of Tyrone is 48 years and his daughter's age is 24
HOPE THIS IS CLEAR TO YOU. IN CASE YOU NEED MORE HELP I AM AVAILABLE FOR ONLINE TUTORING. YOU CAN CONTACT ME ON kapoorjai1@rediffmail.com